# Virtual Reality User Interface Design: Best Practices and Implementation

### v1

@@ Abstract.
Designing effective user interfaces (UIs) for virtual reality (VR) is essential to enhance user immersion, usability, comfort, and accessibility in virtual environments. Despite the growing adoption of VR across domains such as education, healthcare, gaming, and rehabilitation, there is a noticeable lack of unified and comprehensive design guidelines for VR UI design. To address this gap, we conducted a systematic literature review to identify existing best practices and propose complete and unified guidelines for UI development in VR.
@@
Building on these insights, this research proposes a set of best practices to guide the creation of more effective VR interfaces. To demonstrate and validate these practices, we developed a VR application called FlUId that showcases both good and bad UI design principles for direct comparison. A user study was conducted to evaluate the impact of the proposed guidelines. The findings aim to bridge the gap between theory and practice, offering concrete recommendations for VR designers and developers.
Virtual Reality, User Interface Design, Human-Computer Interaction, Usability Guidelines, Immersive Interfaces, VR UX Best Practices, User Study Evaluation.
booktitle: conference: ; 2025; copyright: none
1. Introduction
@@
Virtual Reality (VR) is a computer-generated simulation of a three-dimensional environment that users can interact with and explore in an immersive, often headset-driven, experience. VR originated in the 1960s (Hamdan et al., [n. d.]). It utilizes head-mounted displays (HMDs) to create the illusion of an alternate reality, immersing users in interactive digital environments (Byam, 2019). A well-designed VR experience can enhance user engagement, satisfaction, and immersion (Saranya et al., 2024), but excessive use of the technology may also lead to various health problems and harm personal development (Puspitasari and Lee, 2022). This highlights the importance of thoughtful design, as interface elements and usability directly influence users’ mental models and decisions, with user needs and experiences serving as critical design criteria (Chen et al., 2024).
@@
While gesture-based interface components continue to emerge, traditional user interface (UI) elements remain popular to this day (Yeo et al., 2024). However, for a truly immersive VR experience, it is essential to explore different UI design approaches that support a more natural user interaction (Kharoub et al., 2019). Current methods for designing virtual 3-dimensional (3D) interfaces are typically based on direct adaptations of the desktop metaphor and the related ”window, icon, menu, pointing device” (WIMP) paradigm, resulting in a user experience (UX) gap between physical and virtual reality (Bermejo et al., 2021).
@@
This adaptation is problematic because interaction in 3D virtual environments involves mapping traditional 2-dimensional (2D) input devices to 3D space, which limits how content is perceived when 3D elements are shown as 2D rendered images (Caputo et al., 2019). Unlike conventional 2D interfaces used on desktops and mobile devices, VR environments do not have fixed UI placement boundaries, making traditional UI design principles less directly applicable (Pandey and Sorathia, 2024).
@@
The current lack of established design guidelines for VR interfaces leads to significant limitations, particularly in selecting appropriate interaction modes for immersive experiences (Chen et al., 2024). Existing VR interaction models struggle with user fatigue, motion sickness, and the overall difficulty of designing interfaces that ensure usability and user well-being (Zhou et al., 2023; Khan et al., 2024). There is a lack of established principles for displaying text, with open questions about optimal contrast ratios and other visual parameters affecting readability (Kojić et al., 2020). Additionally, recent studies highlight the challenges in improving menu placement and interaction techniques in VR, emphasizing the need for careful evaluation to enhance usability, accuracy, and user comfort (Byam, 2019; Wang et al., 2021a).
@@
2. Related Work
Designing effective UIs for VR systems is crucial to address immersive environments’ unique challenges and opportunities. With applications ranging from education to mindfulness training and various interaction techniques, understanding how to design these interfaces is key.
2.1. Applications of Virtual Reality
@@
Immersive learning environments are used in education to create engaging and interactive experiences (Sobchyshak et al., 2025). The study (Gonzalez-Argote and Gonzalez-Argote, 2023) proposed a set of best practices for immersive learning design, emphasizing realistic interaction within virtual environments. The work by Xie (Xie, 2023) complemented this by identifying key design points that shape user interface design in educational VR applications, particularly focusing on information clarity, interaction flow, and spatial orientation. The research by Zhou et al. (Zhou et al., 2023) presented a set of 3D user interface design paradigms tailored for VR exhibitions to enhance usability and immersion in virtual environments.
@@
Beyond education, VR is widely adopted in gaming and entertainment. Wang et al. (Wang et al., 2021b) explored design strategies that support collaboration, interactivity, and authenticity in VR games, highlighting how these principles enhance user experience. Works like Larocco (LaRocco, 2020), Alves et al. (Alves et al., 2020), and Çatak et al. (Çatak et al., 2020) analyzed existing industry best practices and manufacturer guidelines, like Oculus and Unity, to propose frameworks that can be generalized across different VR platforms. Kelly et al. (Kelly et al., 2022) presented a VR application designed to support mindfulness practices, emphasizing user comfort, customizable environments, guidance levels, and adaptable interaction modes to foster mindful awareness in safe, controlled settings.
@@
VR interfaces facilitate treatment, recovery, and user well-being in healthcare and rehabilitation (Berrezueta-Guzman et al., 2025). Ramaseri Chandra et al. (Ramaseri Chandra et al., 2022) conducted a systematic survey addressing the causes of cybersickness, a recurring issue in VR environments, and compiled guidelines for improving the virtual experience by minimizing discomfort. Veličković and Milovanović (Veličković and Milovanović, 2021) examined specific design factors in VR rehabilitation, demonstrating that user navigation approaches are the primary cause of cybersickness. Yahaya et al. (Yahaya et al., 2022) further contributed by addressing cybersickness reduction strategies in immersive VR applications, particularly within road safety education for children.
@@
Accessibility and inclusivity are also key themes in current VR research. Zhang et al. (Zhang et al., 2025) derived a set of design guidelines to support inclusive avatar design for people with disabilities, covering avatar appearance, body dynamics, assistive technologies, peripherals, and customization control. Ha et al. (Ha et al., 2022) developed a calibration-free hybrid brain-computer interface system that enables hands-free virtual reality control, particularly targeting users with limited mobility. Ijaz et al. (Ijaz et al., 2022) identified a set of design recommendations for interfaces that accommodate older adults. Kalaiselvi et al. (Kalaiselvi et al., 2024) investigated gesture elicitation in VR for low-vision users, presenting an alternative interaction method that improves immersion, satisfaction, and completion times.
2.2. User Interface Design in Virtual Reality
@@
Current methods for designing virtual 3D interfaces are typically based on direct adaptations of the desktop metaphor and the related ”window, icon, menu, pointing device” (WIMP) paradigm, resulting in a user experience gap between physical and virtual reality (Bermejo et al., 2021). This adaptation is problematic because interaction in 3D virtual environments involves mapping traditional 2D input devices to 3D space, which limits how content is perceived when 3D elements are shown as 2D rendered images (Caputo et al., 2019). VR introduces distinct challenges and factors such as field of view, user orientation, absence of a defined reference plane, and depth-related complexities that differ significantly from flat 2D displays, which need to be considered (Kaur et al., 2019).
@@
Previous research has examined different types of user interfaces, such as graphical user interfaces (GUIs) and natural user interfaces (NUIs), particularly in immersive environments. Regazzoni et al. (Regazzoni et al., 2018) presented guidelines for designing NUIs in VR, focusing on usability, ergonomics, and gesture interaction. The work of Cardona-Reyes et al. (Cardona-Reyes et al., 2020) built on these guidelines to support therapy in children with ADHD. Similarly to NUIs, three-dimensional GUIs lack comprehensive design heuristics, and understanding the cognitive workload of such systems requires considering different variables (Kaur et al., 2020). Kaur and Yammiyavar (Kaur and Yammiyavar, 2017) compared 2D and 3D interface elements to propose guidelines for 3D GUIs.
@@
Research has focused on the design of interface elements like menus, buttons, and keyboards, as well as text readability in VR. Various menu designs, such as Open Palm Menu (Azai et al., 2018a), Tap-tap Menu (Azai et al., 2018b), and an accordion-style menu (Kwon et al., 2017), demonstrate alternative placement options for menus in VR as well as different interaction types. Moreover, the innovative 3D hexa-ring menu introduced by Pandey and Sorathia (Pandey and Sorathia, 2024) showcases a three-dimensional approach that enhances user immersion and ease of interaction, outperforming the traditional 2D design. Kim and Xiong (Kim and Xiong, 2022) introduced pseudo-haptic features for button design that simulate feedback in VR environments, improving user experience and satisfaction. Dudley et al. (Dudley et al., 2019) compared various keyboard typing strategies in VR, concluding that using physical surfaces and two-finger interaction leads to faster typing performance. Dingler et al. (Dingler et al., 2018) investigated factors such as angular height, vergence distance, view box parameters, and text font and background, and identified user preferences for better text readability in VR.
2.3. Interaction and Navigation in Virtual Reality
@@
A crucial aspect of UI design in VR is the selection of effective interaction mechanisms and navigation strategies, which significantly influence user experience. Laine and Suk (Laine and Suk, 2024) investigated how gesture-based interaction affects presence in VR by identifying potential presence factors and proposing a framework for designing and evaluating gesture-based immersive environments. The work by Byam (Byam, 2019) compared three common VR interaction techniques, such as virtual laser pointer, tapping, and d-pad selection, evaluating user performance and preference, and concluded that virtual laser pointer and tapping are the recommended methods for VR selection tasks. Moreover, Kim et al. (Kim et al., 2017) proposed and evaluated a gaze pointer-based user interface for mobile virtual reality, demonstrating how different interface designs, based on visual field range, feedback systems, information transfer methods, and background color, can influence user immersion, presence, convenience, and the risk of VR sickness.
@@
Movement and navigation techniques are also crucial for the overall user experience in VR. Cenydd and Headleand (ap Cenydd and Headleand, 2018) provided an overview of diverse locomotion methods, including teleportation, gamepad locomotion, and other position-tracking methods, and presented design recommendations for movement in virtual reality, particularly within gaming environments. Otaran and Farkhatdinov (Otaran and Farkhatdinov, 2024) explored user preferences for various VR walking interfaces, revealing three preferred VR navigation methods: joystick, point-and-teleport, and walking in place. Byam’s research (Byam, 2019) further explored player navigation strategies, highlighting teleportation as the preferred method over joystick navigation among users (Byam, 2019). Moreover, Bektas et al. (Bektaş et al., 2021) introduced the Limbic Chair, an embodied VR interface that enables independent leg movements, compared to a gamepad. The results showed potential for reducing simulator sickness and enhancing immersion in specific VR scenarios despite slower performance and higher workload.
@@ 3.1. Literature Review
Each of these 97 papers was then examined in detail. During this phase, an additional 42 papers were excluded due to their lack of relevant experimental insight. This process resulted in a final set of 55 papers that formed the analytical basis of this study. The chosen papers were carefully reviewed to identify common themes, design considerations, and established guidelines related to VR UI components and user interaction methods. The full selection process is illustrated in Figure 1.
Figure 1. Literature review process
[…]
This figure shows the process of finding the 55 papers. It illustrates the search methodology, analysis, and results stages.
3.2. Best Practices and Guidelines Development
@@
3.2.1. General VR Design & UI/UX Principles
@@
Alves et al. (Alves et al., 2020) compare manufacturer guidelines, emphasizing depth cues, comfortable placement of elements, and integration of interface elements into virtual environments. LaRocco (LaRocco, 2020) presents the Oculus best practices guide, addressing general user experience considerations including session length management, visual display principles, movement input, and more. Regazzoni et al. (Regazzoni et al., 2018) present natural user interface design guidelines focusing on dialogue principles, ergonomic considerations, and gesture-based interactions for VR environments.
Hari Chandana et al. (Chandana et al., 2023) suggest user experience design best practices in VR, emphasizing the importance of understanding user needs, rapid prototyping, precise wayfinding mechanisms, and intuitive interaction design. Lima et al. (Lima et al., 2024) focus on ergonomics and user experience in VR interface design, highlighting the need to move beyond 2D design principles, design elements in an intuitive way, and incorporate storytelling techniques. Khan et al. (Khan et al., 2024) provide human factors and ergonomic design guidelines in VR, covering performance requirements, accessibility considerations, and visual fatigue mitigation strategies. Kaur and Yammiyavar (Kaur and Yammiyavar, 2017) offer insights into adaptive interface design, incorporating feedback and cognitive load management.
Shoikova and Peshev (Shoikova and Peshev, 2017) establish best practices for VR design, emphasizing the importance of creating believable environments, stimulating users’ senses, and designing avatar-centric interfaces that account for personal space and multitasking. Wall (Wall, 2021) presents empirical findings on VR menu interaction and design, providing recommendations for visualization, haptic feedback, ray-casting, and pre-attentive processing. Yeo et al. (Yeo et al., 2024) offer design recommendations for aligning input device dimensionality with virtual tasks, limiting unnecessary motion, and highlighting the importance of context-dependent interaction.
3.2.2. VR Game Design & Development

### v2

@@ Abstract
Designing effective user interfaces (UIs) for virtual reality (VR) is essential to enhance user immersion, usability, comfort, and accessibility in virtual environments. Despite the growing adoption of VR across domains, there is a noticeable lack of unified and comprehensive design guidelines for VR UI design. To address this gap, we conducted a systematic literature review to identify existing best practices and propose 28 unified guidelines for UI development in VR.
@@
Building on these insights, this research proposes a framework to guide the creation of more effective VR interfaces. To demonstrate and validate these practices, we developed a VR application called FlUId and an interactive Web Tool that serves as a guideline explorer and project planning resource for developers. A user study was conducted to evaluate the impact of the proposed guidelines. The findings aim to bridge the gap between theory and practice, offering concrete recommendations and digital tools for VR designers and developers.
keywords:
[…]
1 Introduction
@@
Virtual Reality (VR) is a computer-generated simulation of a three-dimensional environment that users can interact with and explore in an immersive, often headset-driven, experience. VR originated in the 1960s [hamdanindustry]. It utilizes head-mounted displays (HMDs) to create the illusion of an alternate reality, immersing users in interactive digital environments [byam2019best]. A well-designed VR experience can enhance user engagement, satisfaction, and immersion [saranya2024development], but excessive use of the technology may also lead to various health problems and harm personal development [puspitasari2022review]. This highlights the importance of thoughtful design, as interface elements and usability directly influence users’ mental models and decisions, with user needs and experiences serving as critical design criteria [chen2024survey].
@@
While gesture-based interface components continue to emerge, traditional user interface (UI) elements remain popular to this day [yeo2024entering]. However, for a truly immersive VR experience, it is essential to explore different UI design approaches that support a more natural user interaction [kharoub20193d]. Current methods for designing virtual 3-dimensional (3D) interfaces are typically based on direct adaptations of the desktop metaphor and the related "window, icon, menu, pointing device" (WIMP) paradigm, resulting in a user experience (UX) gap between physical and virtual reality [bermejo2021exploring].
@@
This adaptation is problematic because interaction in 3D virtual environments involves mapping traditional 2-dimensional (2D) input devices to 3D space, which limits how content is perceived when 3D elements are shown as 2D rendered images [caputo2019gestural]. Unlike conventional 2D interfaces used on desktops and mobile devices, VR environments do not have fixed UI placement boundaries, making traditional UI design principles less directly applicable [pandey2024pilot].
@@
The current lack of established design guidelines for VR interfaces leads to significant limitations, particularly in selecting appropriate interaction modes for immersive experiences [chen2024survey]. Existing VR interaction models struggle with user fatigue, motion sickness, and the overall difficulty of designing interfaces that ensure usability and user well-being [zhou2023design, khanvirtual]. There is a lack of established principles for displaying text, with open questions about optimal contrast ratios and other visual parameters affecting readability [kojic2020user]. Additionally, recent studies highlight the challenges in improving menu placement and interaction techniques in VR, emphasizing the need for careful evaluation to enhance usability, accuracy, and user comfort [byam2019best, wang2021experimental].
@@
1.1 Applications of Virtual Reality
@@
Immersive learning environments are used in education to create engaging and interactive experiences [sobchyshak2025]. The study [gonzalez202310] proposed a set of best practices for immersive learning design, emphasizing realistic interaction within virtual environments. The work by Xie [Xie] complemented this by identifying key design points that shape user interface design in educational VR applications, particularly focusing on information clarity, interaction flow, and spatial orientation. The research by Zhou et al. [zhou2023design] presented a set of 3D user interface design paradigms tailored for VR exhibitions to enhance usability and immersion in virtual environments.
@@
Beyond education, VR is widely adopted in gaming and entertainment. Wang et al. [wang2021authenticity] explored design strategies that support collaboration, interactivity, and authenticity in VR games, highlighting how these principles enhance user experience. Works like Larocco [larocco2020developing], Alves et al. [alves2020evaluation], and Çatak et al. [ccatak2020guideline] analyzed existing industry best practices and manufacturer guidelines, like Oculus and Unity, to propose frameworks that can be generalized across different VR platforms. Kelly et al. [kelly2022design] presented a VR application designed to support mindfulness practices, emphasizing user comfort, customizable environments, guidance levels, and adaptable interaction modes to foster mindful awareness in safe, controlled settings.
@@
VR interfaces facilitate treatment, recovery, and user well-being in healthcare and rehabilitation [berrezuetaguzman2025]. Ramaseri Chandra et al. [ramaseri2022systematic] conducted a systematic survey addressing the causes of cybersickness, a recurring issue in VR environments, and compiled guidelines for improving the virtual experience by minimizing discomfort. Veličković and Milovanović [velivckovic2021improvement] examined specific design factors in VR rehabilitation, demonstrating that user navigation approaches are the primary cause of cybersickness. Yahaya et al. [yahaya2022comparative] further contributed by addressing cybersickness reduction strategies in immersive VR applications, particularly within road safety education for children.
@@
Accessibility and inclusivity are also key themes in current VR research. Zhang et al. [zhang2025inclusive] derived a set of design guidelines to support inclusive avatar design for people with disabilities, covering avatar appearance, body dynamics, assistive technologies, peripherals, and customization control. Ha et al. [ha2022novel] developed a calibration-free hybrid brain-computer interface system that enables hands-free virtual reality control, particularly targeting users with limited mobility. Ijaz et al. [ijaz2022design] identified a set of design recommendations for interfaces that accommodate older adults. Kalaiselvi et al. [kalaiselvi] investigated gesture elicitation in VR for low-vision users, presenting an alternative interaction method that improves immersion, satisfaction, and completion times.
1.2 User Interface Design in Virtual Reality
@@
Current methods for designing virtual 3D interfaces are typically based on direct adaptations of the desktop metaphor and the related "window, icon, menu, pointing device" (WIMP) paradigm, resulting in a user experience gap between physical and virtual reality [bermejo2021exploring]. This adaptation is problematic because interaction in 3D virtual environments involves mapping traditional 2D input devices to 3D space, which limits how content is perceived when 3D elements are shown as 2D rendered images [caputo2019gestural]. VR introduces distinct challenges and factors such as field of view, user orientation, absence of a defined reference plane, and depth-related complexities that differ significantly from flat 2D displays, which need to be considered [kaur2019exploring].
@@
Previous research has examined different types of user interfaces, such as graphical user interfaces (GUIs) and natural user interfaces (NUIs), particularly in immersive environments. Regazzoni et al. [regazzoni2018virtual] presented guidelines for designing NUIs in VR, focusing on usability, ergonomics, and gesture interaction. The work of Cardona-Reyes et al. [cardona2020natural] built on these guidelines to support therapy in children with ADHD. Similarly to NUIs, three-dimensional GUIs lack comprehensive design heuristics, and understanding the cognitive workload of such systems requires considering different variables [kaur2020cognitive]. Kaur and Yammiyavar [kaur2017comparative] compared 2D and 3D interface elements to propose guidelines for 3D GUIs.
@@
Research has focused on the design of interface elements like menus, buttons, and keyboards, as well as text readability in VR. Various menu designs, such as Open Palm Menu [azai2018open], Tap-tap Menu [azai2018tap], and an accordion-style menu [kwon2017spatial], demonstrate alternative placement options for menus in VR as well as different interaction types. Moreover, the innovative 3D hexa-ring menu introduced by Pandey and Sorathia [pandey2024pilot] showcases a three-dimensional approach that enhances user immersion and ease of interaction, outperforming the traditional 2D design. Kim and Xiong [kim2022pseudo] introduced pseudo-haptic features for button design that simulate feedback in VR environments, improving user experience and satisfaction. Dudley et al. [dudley2019performance] compared various keyboard typing strategies in VR, concluding that using physical surfaces and two-finger interaction leads to faster typing performance. Dingler et al. [dingler2018vr] investigated factors such as angular height, vergence distance, view box parameters, and text font and background, and identified user preferences for better text readability in VR.
1.3 Interaction and Navigation in Virtual Reality
@@
A crucial aspect of UI design in VR is the selection of effective interaction mechanisms and navigation strategies, which significantly influence user experience. Laine and Suk [laine2024investigating] investigated how gesture-based interaction affects presence in VR by identifying potential presence factors and proposing a framework for designing and evaluating gesture-based immersive environments. The work by Byam [byam2019best] compared three common VR interaction techniques, such as virtual laser pointer, tapping, and d-pad selection, evaluating user performance and preference, and concluded that virtual laser pointer and tapping are the recommended methods for VR selection tasks. Moreover, Kim et al. [kim2017study] proposed and evaluated a gaze pointer-based user interface for mobile virtual reality, demonstrating how different interface designs, based on visual field range, feedback systems, information transfer methods, and background color, can influence user immersion, presence, convenience, and the risk of VR sickness.
@@
Movement and navigation techniques are also crucial for the overall user experience in VR. Cenydd and Headleand [ap2018movement] provided an overview of diverse locomotion methods, including teleportation, gamepad locomotion, and other position-tracking methods, and presented design recommendations for movement in virtual reality, particularly within gaming environments. Otaran and Farkhatdinov [otaran2024exploring] explored user preferences for various VR walking interfaces, revealing three preferred VR navigation methods: joystick, point-and-teleport, and walking in place. Byam’s research [byam2019best] further explored player navigation strategies, highlighting teleportation as the preferred method over joystick navigation among users [byam2019best]. Moreover, Bektas et al. [bektacs2021systematic] introduced the Limbic Chair, an embodied VR interface that enables independent leg movements, compared to a gamepad. The results showed potential for reducing simulator sickness and enhancing immersion in specific VR scenarios despite slower performance and higher workload.
@@ 2.1 Literature Review
Each of these 97 papers was then examined in detail. During this phase, an additional 42 papers were excluded due to their lack of relevant experimental insight. This process resulted in a final set of 55 papers that formed the analytical basis of this study. The chosen papers were carefully reviewed to identify common themes, design considerations, and established guidelines related to VR UI components and user interaction methods. The full selection process is illustrated in Figure 2.1.
Figure 1: Literature review process
2.2 Best Practices and Guidelines Development
@@
2.2.1 General VR Design & UI/UX Principles
@@
Alves et al. [alves2020evaluation] compare manufacturer guidelines, emphasizing depth cues, comfortable placement of elements, and integration of interface elements into virtual environments. LaRocco [larocco2020developing] presents the Oculus best practices guide, addressing general user experience considerations including session length management, visual display principles, movement input, and more. Regazzoni et al. [regazzoni2018virtual] present natural user interface design guidelines focusing on dialogue principles, ergonomic considerations, and gesture-based interactions for VR environments.
Hari Chandana et al. [chandana2023exploring] suggest user experience design best practices in VR, emphasizing the importance of understanding user needs, rapid prototyping, precise wayfinding mechanisms, and intuitive interaction design. Lima et al. [lima2024construction] focus on ergonomics and user experience in VR interface design, highlighting the need to move beyond 2D design principles, design elements in an intuitive way, and incorporate storytelling techniques. Khan et al. [khanvirtual] provide human factors and ergonomic design guidelines in VR, covering performance requirements, accessibility considerations, and visual fatigue mitigation strategies. Kaur and Yammiyavar [kaur2017comparative] offer insights into adaptive interface design, incorporating feedback and cognitive load management.
Shoikova and Peshev [shoikova2017best] establish best practices for VR design, emphasizing the importance of creating believable environments, stimulating users’ senses, and designing avatar-centric interfaces that account for personal space and multitasking. Wall [wall2021empirical] presents empirical findings on VR menu interaction and design, providing recommendations for visualization, haptic feedback, ray-casting, and pre-attentive processing. Yeo et al. [yeo2024entering] offer design recommendations for aligning input device dimensionality with virtual tasks, limiting unnecessary motion, and highlighting the importance of context-dependent interaction.
2.2.2 VR Game Design & Development